OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (anonymous):

Hint: Expand sin3x by writing it as : sin(x+2x)

myininaya (myininaya):

Recall 3x=2x+x So sin(3x)=sin(x+2x) Use the formula sin(u+v)=sin(u)cos(v)+sin(v)cos(u) to expand sin(x+2x)

OpenStudy (anonymous):

i was thinking that but i wasnt sure

OpenStudy (anonymous):

now all the options i am given have exponents in them how is that possible

OpenStudy (anonymous):

That will be the case. sin2x is 2sinxcosx.

OpenStudy (anonymous):

i dont understand...